An acquirer with a P/E ratio of 13 and earnings of $1.42 seeks to take over another target firm with value $12.57 and P/E ratio 15. What is the new merged firm's P/E?
Answer #1

Acquirer (A) P/E ratio = 13

Earning of Acquirer(A) (EPS-A) = $1.42

Target firm (B) value (Market Price per share(MPS-B)) = $12.57

Target firm(B) P/E ratio = 15

We have following relationship and formula -

MPS- EPS* P/E

Thus,

MPS(A) = 1.42 * 13

MPS(A) = 18.46

Value of Acquirer per share is $18.46

And,

MPS(B) = EPS(B) * P/E(B)

EPS(RMPS B PE(B) EPSE

12.57 EPS(B)- 15

EPS(B) = 0.838

Earning of Target firm is $0.838

MPS(A) +MPS(B) EPS(A) EPS(B) Post -mergerE

8.4612.57 1.42 + 0.838 Post merger P/E-

31.03 58 Post -merger P/E

Post-merger P/E = 13.7422497785

Thus,

New merged firm's P/E 13.74

Hope this will help, please do comment if you need any further explanation. Your feedback would be appreciated.